OpenStudy (anonymous):

can anyone please show me the steps to solve this problem :) 3=x*3^4 but in expotenetial form :)

OpenStudy (anonymous):

\[3=x\times 3^4\]?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Sorry, but you need to clarify a bit here.. Is the equation really this: \(3=x\cdot3^4\) Or is it this? \(3 = (x\cdot3)^4\)

OpenStudy (anonymous):

would u like to see the whole question ?

OpenStudy (anonymous):

Yes, that would be helpful.

OpenStudy (anonymous):

if it is the one i posted, divide both sides by \[3^4\] then simplify then take the cube root

OpenStudy (anonymous):

divive both by 29 ?

OpenStudy (anonymous):

i got 0.103=x

OpenStudy (anonymous):

well actually 3^4=81 but yes that is the idea

OpenStudy (anonymous):

Oh, ok. So you have: \[3 = x\cdot 3^4\]\[\implies x = \frac{3}{3^4} = \frac{1}{3^3} = \frac{1}{27}\]

OpenStudy (anonymous):

oh dear forget decimals

OpenStudy (anonymous):

what polpak said. hello polpak!

OpenStudy (anonymous):

Decimals are like calculators. To be avoided as often as possible.

OpenStudy (anonymous):

oh sorry i did 3^3

OpenStudy (anonymous):

oh so how do you put 1/27 in expotential form ? thats what i need help with ; i get the whole precess now :) thanks much it was very helpful ;' so how to put in expotentail form?

OpenStudy (anonymous):

Oh. exponential form.. \[\frac{1}{3^3} = 3^{-3}\]

OpenStudy (anonymous):

Just have to turn the power in the denominator into a factor with a negative exponent.

OpenStudy (anonymous):

ohhh! :D got it ! thanks soo much :) and everything you said i wrote in noters :) thanks again .